OpenStudy (anonymous):

The number 2005 has 2 prime factors. How many positive factors exist for 2005^2005?

OpenStudy (anonymous):

2005^2005 = 2005*2005*2005...2005 times And 2005 has 2 prime factors Thus 2005^2005 will have 2*2005 prime factors! = 4010 factors!

OpenStudy (anonymous):

Thanks